Assuming a private closed economy whereby the marginal propensity to
consume is 0.9 and investment spending decreases by R1000 billion. What will
be the change on equilibrium GDP?
1
Expert's answer
2020-04-21T18:44:17-0400

The investment multiplier is computed as:



"\color{red}{\dfrac{\Delta Y}{\Delta I} = \frac{1}{1-MPC}}"

Where MPC is the marginal propensity to consume.


Given that MPC = 0.9, the investment multiplier will be equal to:


"\dfrac{\Delta Y}{\Delta I} =\dfrac{1}{1-0.9}=10"

This means that an increase in investment spending by 1 unit will increase the equilibrium GDP by 10 units.


Thus, increasing the investment spending by R1000 billion will increase the equilibrium GDP by:



"\dfrac{\Delta Y}{\rm R1,000} = 10"


"\Delta Y= 10\times \rm R1,000\; billion=\boxed{\color{blue}{\rm R10,000\; billion}}"